A company predicts that it will sell y units of its product if it charges x dollars for it using the formula y=300-2.5x. In whole dollars, how much should the company charge for the product if it wants to sell 175 units.
Answer by lwsshak3(11628)

A company predicts that it will sell y units of its product if it charges x dollars for it using the formula y=300-2.5x. In whole dollars, how much should the company charge for the product if it wants to sell 175 units.
**
This equation is for a straight line of the standard form: y=mx+b, m=slope, b=y-intercept
175=300-2.5x
y=-2.5x+300
175=-2.5x+300
2.5x=300-175
2.5x=125
x=125/2.5=50
ans:
The company charge for the product should be $50